OpenStudy (anonymous):

432, 455, 280, 216, 353, 387, 119, 293, 241, 460, 389, 121, 142 Find the mean of the data set. Find the median of the data set. Find the first and third quartiles of the data set

OpenStudy (anonymous):

The median is when you add the biggest number and the lowest number and divide in this case 460+142=602/2= 301

OpenStudy (anonymous):

i really need the 3rd part of this question

OpenStudy (anonymous):

first find least number to greatest number 119, 121, 142, 216, 241, 280, 293, 353, 387, 389, 432, 455, 460 the median is 293. How do i get the answer. well you have to find which number is medium. The mean is 299. How do i get the answer. You have to add all together 119 + 121 + 142 + 216 + 241 + 280 + 293 + 353 + 387 + 389 + 432 + 455 + 460 = 3888 119, 121, 142, 216, 241, 280, 293, 353, 387, 389, 432, 455, 460 1 2 3 4 5 6 7 8 9 10 11 12 13 3888/13 = 299 and sorry i don't know the third one...